The letters of the word Mississippi are placed in a hat. What is the probability that the first two letters drawn are both i if the first letter drawn is not replaced?
Answer by Fombitz(32388)

MISSISSIPPI
1 M's
4 I's
4 S's
2 P's
11 total
P(first I)=4/11
P(second I)=3/10
P(2 I's)=(4/11)(3/10)=6/55